Table 2.2: U.S. 2014-2015 Domestic Income (S billions) 2014 9,2649,704 7,487 1,777 1,210,238 2015 Compensation of employees, paid Wages and salaries Supplements to wages and salaries 7,866 1,838 Business taxes 57 4,489 4,509 20 2,745 57 4,575 4,593 19 2,831 Business subsidieS Net operating surplus Private enterprises Surplus of government enterprises Depreciation of fixed capital (Source: Bureau of Economic Analysis)

How to calculate 2014, 2015 GDP? Also the net domestic product of 2014, 2015.

Answers should be: (2014 GDP: $17,651, 2015 GDP: $18,290
, 2014 dp:
a. $14,906, 2015 dp: e. $15,460 )
Answer #1

Solution:

National Income or Net Domestic Product = Compensation of employees (i.e. wages and salaries + supplements to wages and salaries) + Net Operating Surplus (i.e. private enterprises + government enterprises) +(Business taxes - Business Subsidies)

And Gross domestic product GDP = NDP +Depreciation or capital consumption allowance + adjustments

Now NDP for 2014 = 9264+4489+1210-57 = $14,906 billion

NDP for 2015 = 9704+4575+1238-57 = $15460 billion

GDP for 2014 = 14906 + 2745 = $17,651 billion

GDP for 2015 = 15460 + 2831 = $18,291 billion